1.      Given the following cost function: C(x) = 3x+2/x+4 ,

a)     Find the MC function                                                                                                        

b) What is the marginal cost when  x=20


1
Expert's answer
2021-09-03T08:09:37-0400

a) The MC function is:

"MC = C'(x) = 3 - 2\/x^2."

b) If x = 20, then "MC = 3 - 2\/400 = 2.995."